OpenStudy (anonymous):

What is the solution to y over two + 5 = 9? 28 7 8 2

OpenStudy (charlotte123):

y/2 + 5 = 9 - 5 - 5 y/2 = 4 * 2 * 2 y = 8

OpenStudy (anonymous):

the solution exists!

OpenStudy (charlotte123):

TA DA

OpenStudy (charlotte123):

BEHOLD THE SOLUTION

OpenStudy (charlotte123):

*Insert sparkles here*

OpenStudy (anonymous):